A winter coat manufacturer is deciding whether to manufacture a puffer coat or a parka. For the puffer coat the cost of producing each item is $50.55 and they will sell each item to a reseller for $167.48. For the parka the cost of producing each item is $65.66 but they will sell each item to a reseller for $200.98.


What is the contribution rate for manufacturing the puffer coats? Your answer should be accurate to at least two decimal places.

What is the contribution rate for manufacturing the parkas? Your answer should be accurate to at least two decimal places.

Which type of coat has the better contribution rate?

Pleas
Answer #1

Contribution Margin = Sales Price per Unit - Variable Cost per Unit

For Puffer Coat

Contribution Margin per Unit = 167.48 - 50.55

Contribution Margin per Unit = $ 116.93

For Parka Coats

Contribution Margin per Unit = 200.98 - 65.66

Contribution Margin per Unit = $ 135.32

Contribution Rate = Contribution Margin per Unit / Sales Price per Unit * 100

For Puffer Costs

Contribution Rate = 116.93 / 167.48 * 100

Contribution Rate = 69.81%

For Parka Coats

Contribution Margin Rate = 135.32 / 200.98 * 100

Contribution Margin Rate = 67.33%

Puffer Costs has better Contribution Rate.

Notes

Cost of Production is treated as a Variable Cost.